# Discovering Hidden Visual Concepts Beyond Linguistic Input in Infant Learning

### v1

@@ 4.2 Results
Figure 7: In- and out-of-vocabulary class performance across $n$-way settings. While CVCL has limited training vocabulary (in-vocab), our NeuronClassifier framework successfully uncovers hidden visual concepts, enabling out-of-vocabulary classification without additional training. The “Vanilla” represents the classification outlined in Figure 2. The baseline (black) indicates chance-level performance. Among these $n$-way settings, our method maintains the relative performance over others for both in-vocab and out-of-vocab words, supporting the presence of beyond-vocabulary potential in infant models.
We evaluate the models under various $n$-way classification setups ($n=2,4,8,16,32$). Figure 7 illustrates the performance trends for out-of-vocabulary class classification accuracy using our proposed method, NeuronClassifier.
Our method not only enables out-of-vocabulary classification but also significantly improves in-vocab performance compared to previous results [38], further supporting the presence of beyond-vocabulary potential in infant models. However, due to limited class coverage, the CVCL model with our method can classify a maximum of 31 classes, resulting in missing values at $n=32$. These findings support our hypothesis that the infant model has acquired visual concepts beyond its initial vocabulary. These results show that leveraging the model’s internal representations for classification that go beyond its vocabulary is not only feasible but also robust across different $n$-way settings. These findings support our hypothesis that the infant model has acquired visual concepts beyond its initial vocabulary.
@@ 5 Layer-wise Representation Analysis
Figure 8: CKA layer-wise similarity between CVCL and ResNeXt50. Using ImageNet validation set as $\mathcal{D}_{\textrm{probe}}$, CVCL demonstrates similar capabilities to CLIP-RN50 and ResNeXt50 in the shallow layers (lower-level features), but diverges significantly in the final layer (higher-level features).
While the previous section demonstrated that the model trained on infant data learned visual concepts beyond parental speech, its classification performance still lagged behind models trained on large-scale datasets like ImageNet. To explore how the representations captured by infant models compare to those of large-scale pre-trained models, we perform a layer-wise representation analysis using Centered Kernel Alignment (CKA) [23]. This method measures the similarity between the representations of the CVCL-ResNeXt50 model, an ImageNet-pretrained ResNeXt, and CLIP-ResNet50. Additionally, we apply neuron labeling techniques from a layer-wise perspective to identify unique visual concepts discovered at each layer. We found that the large-scale pre-trained models have much more diversity of high-level visual concepts in the last few layers, which aligns with the low similarity observed in these layers on CKA when compared to CVCL. This indicates that CVCL lacks the development of a certain level of complexity in visual concepts, providing insights into how the infant model and large-scale pre-trained models diverge or align at different levels of abstraction.
Centered Kernel Alignment (CKA).
@@
CKA is an effective metric for comparing the similarity between the layer-wise representations of different neural networks. For two sets of activations, $\mathbf{X}\in\mathbb{R}^{n\times p}$ and $\mathbf{Y}\in\mathbb{R}^{n\times q}$, from corresponding layers of two models, where $n$ is the number of examples and $p$ and $q$ are the feature dimensions (i.e., the number of neurons in each layer), the linear CKA is defined as:
@@
Layer-wise similarity.
To obtain activations $\mathbf{X}$ and $\mathbf{Y}$, we used the ImageNet validation set [10] as input for the networks. We computed the CKA similarity between the infant model (CVCL), the ImageNet-pretrained ResNeXt50, and CLIP-ResNet50. The results are presented as matrices in Figure 8.
Our layer-wise analysis using CKA highlights the differences in how CVCL captures representations compared to well-pretrained models. Layers closer to the input exhibit higher similarity across models, while deeper layers show greater divergence. This pattern aligns with findings in model training, where shallow layers tend to converge early on low-complexity features, and deeper layers progressively specialize to capture higher-level concepts [13, 7].
Despite being trained on limited, noisy, and sparse data, CVCL successfully develops lower-level representations that are comparable to those in large-scale pre-trained models. However, the divergence in deeper layers reflects the constraints imposed by its dataset, indicating a lack of the more diverse high-level representations typically observed in models trained on large-scale datasets.
Unique visual concept neurons identification.
Figure 9: Unique concepts discovered between infant and ImageNet model. Neurons in deeper model layers capture increasingly complex concepts. Early layers primarily detect lower-level features like color and texture, while higher-level concepts such as objects and scenes emerge in deeper layers. The CVCL model exhibits fewer unique detectors than the ImageNet model, especially for high-level visual concepts (e.g., objects and scenes).
To further investigate the unique characteristics of each layer, we apply network dissection to identify neurons that are aligned with specific visual concepts and are unique. Essentially, this is an extension of the neuron labeling process, where the Broden[2] dataset provides category labels for each visual concept neuron. We counted the number of unique visual concepts discovered in each category and perform layer-wise comparisons to gain a broader view of the differences between models. In Figure 9, we visualize the number of unique detectors across layers for each model. The results show that early layers in all models predominantly detect low-level features like color, with minimal differences between models. However, as we move to deeper layers, high-level concepts such as objects and scenes become more prominent, and the disparities between models become clearer. The CVCL model exhibits fewer unique detectors in these high-level categories compared to ImageNet model, reflecting its limited exposure to diverse, large-scale datasets.
6 Conclusion
In this paper, we explored whether an infant-like model (CVCL), trained on infant egocentric video frames and linguistic inputs, can acquire broader visual concepts extending beyond its initial training vocabulary. By introducing NeuronClassifier, a training-free framework to discover and leverage visual concepts hidden in representations, we unlocked the CVCL visual encoder’s ability to recognize out-of-vocabulary concepts, establishing its potential as a strong classifier. Our findings also reveal that while CVCL’s representations capture low-level features similar to those in large-scale pre-trained models, they diverge significantly in higher-level representations, contributing to the observed performance differences.
Overall, our approach bridges cognitive science and computer vision, providing insights into how infant-like models develop visual concepts that precede linguistic inputs, aligning with the natural way infants explore the world through sight.
Appendix A Neuron-wise Analysis
We present additional examples illustrating how the infant model classifies images using visual-concept neurons. Furthermore, we provide complete results of Age of Acquisition (AoA) to quantify the cognitive level of visual concepts for both in-vocabulary and out-of-vocabulary cases.
A.1 Neuron-based Classification Examples
A.2 Age-of-Acquisition Ratings
A.2.1 Details
Age-of-Acquisition (AoA) ratings, as defined by Kuperman et al. [25], estimate the age at which a person learns a word. These ratings were obtained via crowdsourcing using 30,121 English content words, organized into frequency-matched lists based on the SUBTLEX-US corpus [4]. Each list included calibrator and control words for validation.
The AoA ratings strongly correlated with prior norms ($r=0.93$ with Cortese and Khanna [9], $r=0.86$ with the Bristol norms [34]), confirming their reliability for studying vocabulary development.
A.2.2 Detailed AoA Results
Here we present visual concepts that originally from Konkle object dataset[22] class names, by applying neuron labeling techniques, we found many visual-concept neurons with corresponding class inside vision encoder’s hidden representation. For founded classes, we investigate their AoA values to prove the alignment between computational model and infant cognition. The out-of-vocab class in Table 3 tend to have higher AoA values than in-vocab in Table 2, representing the early development of infant visual system that extend beyond linguistic inputs.
Vocab (Col 1) AoA (Col 1) Vocab (Col 2) AoA (Col 2)
bike 2.9 camera 4.78
stamp 2.94 umbrella 4.79
microwave 3.23 bagel 4.79
pen 3.33 desk 5.00
knife 3.37 hat 5.11
broom 3.43 cookie 5.50
scissors 4.05 stool 5.56
hairbrush 4.15 necklace 5.61
button 4.15 sofa 5.63
pizza 4.26 fan 5.68
kayak 4.42 chair 6.00
bucket 4.5 ball 6.21
clock 4.5 sandwich 6.33
apple 4.67 pants 7.67
tricycle 4.7 socks (sock) 8.80
bowl 8.90
Vocab (Col 1) AoA (Col 1) Vocab (Col 2) AoA (Col 2)
frisbee 5.50 domino 6.17
hammer 5.42 calculator 8.22
chessboard 8.37 rosary 8.21
necktie 6.63 licenseplate (license) 8.70
hanger 6.78 ceilingfan (fan) 5.63
muffins (muffin) 5.11 telescope 6.95
keyboard 9.32 golfball (golf) 7.16
toyhorse (horse) 4.15 yarn 6.61
cushion 5.53 dresser 4.28
rug 4.61 grill 6.53
helmet 5.71 cupsaucer (saucer) 8.44
stapler 5.83 pokercard (poker) 9.10
dollhouse 4.86 recordplayer 6.37
axe 6.11 speakers (speaker) 6.11
tent 5.16 roadsign (sign) 4.32
watergun (gun) 5.58 sippycup (cup) 3.57
binoculars 6.79 lawnmower 6.11
toyrabbit (rabbit) 3.94 trunk 8.30
suitcase 8.22 ringbinder (binder) 10.42
compass 8.44 powerstrip 12.01
bowtie 7.94 doorknob 4.70
lantern 8.55 mask 4.80
bathsuit (bathrobe) 7.90 collar 6.56
dumbbell 7.56 seashell 7.06
pitcher 6.42

### v4

@@ 4.2 Results
We evaluate the models under various $n$-way classification setups. Figure 8 illustrates the performance trends for in- and out-of-vocabulary class classification accuracy applying our NeuronClassifier.
Our method not only enables out-of-vocabulary classification but also significantly improves in-vocab performance compared to previous results [43], further supporting the presence of beyond-vocabulary potential in infant models. However, due to limited class coverage, CVCL with our method can classify a maximum of 31 classes (see Appendix A.3). These findings support our hypothesis that the infant model has acquired visual concepts beyond its initial vocabulary. These results show that leveraging the model’s internal representations for classification that go beyond its vocabulary is not only feasible but also robust across different $n$-way settings. These findings support our hypothesis that the infant model has acquired visual concepts beyond its initial vocabulary.
@@
Figure 8: In- and out-of-vocabulary class performance across $n$-way settings using NeuronClassifier. (a) For out-of-vocabulary classification (left), our method enables classification without additional training. The infant model CVCL ( ) maintains robust performance as $n$ increases. (b) For in-vocabulary classification (right), “ours” (NeuronClassifier) enhances CVCL recognition ability compared to the “vanilla” setting. However, CLIP ( ) declines, as it relies solely on neurons without text encoder input. Random ones serve as lower bounds.
@@ 5 Layer-wise Representation Analysis
How does the representation learned from infant data differ from other representations widely used in the computer vision community? To explore this, we perform a layer-wise representation analysis using Centered Kernel Alignment (CKA) [26]. We compute the similarity between the representations of the infant model(CVCL), ImageNet-pretrained, and CLIP. Additionally, we apply neuron labeling techniques from a layer-wise perspective to identify unique visual concepts discovered at each layer between ImageNet and infant models.
Layer similarity analysis.
CKA [26] applies HSIC [20] (see Appendix B) over a set of images to provide layer-wise similarity scores between the representations of different neural networks. A higher CKA score indicates more similar representations between two models at the given layer. We use the ImageNet validation set [11] as input for the networks, and compute the CKA similarity between the infant model (CVCL), the ImageNet-pretrained ResNeXt50, and CLIP-ResNet50. The results are presented as matrices in Figure 9. The lower layers of CVCL exhibit greater similarity to larger-scale models than its deeper layers. In larger-scale models, shallow layers are known to capture low-complexity features, while deeper layers progressively specialize in capturing higher-level concepts [15, 7]. Therefore, CVCL – the model trained on infant data – successfully develops lower-level representations comparable to those in common pre-trained models. However, the divergence in deeper layers suggests a lack of the diverse higher-level representations typically observed in models trained on common datasets.
@@
Figure 9: CKA layer-wise similarity between CVCL and common models. Using the ImageNet validation set as input, CVCL (y-axis) exhibits similarity to CLIP-RN50 (x-axis, left) and ResNeXt50 (x-axis, right) in the shallow layers (lower-level features) but diverges significantly in the final layer (higher-level features). Notably, Layer 4 of CVCL shows very low similarity to all layers of both common models.
Neuron-based analysis.
To investigate the characteristics of each layer, we apply network dissection to identify neurons that are aligned with specific visual concepts. Essentially, this is an extension of the neuron labeling process, where the Broden [2] dataset provides category labels for each visual concept neuron. We count the number of unique visual concepts discovered in each category and perform layer-wise comparisons to gain a broader view of the differences between models trained from ImageNet and infant data. In Figure 10, we visualize the number of unique visual concept neurons across layers for each model. The results show that early layers in both models predominantly have neurons of low-level features like color, with minimal differences between models. However, as we move to deeper layers, higher-level concepts such as objects and scenes become more prominent, and the disparities between models become clearer. CVCL exhibits fewer unique visual concepts in these higher-level categories compared to ImageNet model. This finding aligns with the layer similarity analysis.
Figure 10: Number of visual concepts in ImageNet and infant models. Concepts are categorized using Broden dataset [2]. Neurons in deeper model layers capture increasingly complex concepts. Early layers primarily detect lower-level features like color and texture, while higher-level concepts such as objects and scenes emerge in deeper layers. CVCL exhibits fewer visual concepts than the ImageNet model, especially for higher-level visual concepts (e.g., objects and scenes).
6 Conclusion
In this paper, we explored whether an infant model (CVCL), trained on infant egocentric video frames and linguistic inputs, can acquire broader visual concepts extending beyond its initial training vocabulary. By introducing NeuronClassifier, a training-free framework to discover and leverage visual concepts hidden in representations, we unlocked the CVCL visual encoder’s ability to recognize out-of-vocabulary concepts, establishing its potential as a strong classifier. Our findings also reveal that while CVCL, trained on a unique infant dataset with limited exposure to diverse scenes, it representations capture low-level features similar to those in common pre-trained models, they diverge significantly in higher-level representations, contributing to the observed performance differences.
Overall, our approach bridges cognitive science and computer vision, providing insights into how infant models develop visual concepts that precede linguistic inputs, aligning with the natural way infants explore the world through sight.
Limitations and future work.
Our study did not analyze the infant training data, as we were unable to access it due to limited access controls (see Appendix C). Instead, we analyze models trained on infant egocentric data, finding developmental alignment with cognitive studies. However, we did not extend this analysis to adult egocentric data. While the study focuses on infant data for developmental process, the framework can be applied to adult models, which remains a direction for future research.
Acknowledgements
This work was partially supported by the National Research Foundation Singapore Competitive Research Program (award number CRP29-2022-0003). We thank Wai Keen Vong for invaluable discussion and CVCL’s pretrained weights. We thank Jingyi Lin for figure discussions. We thank the anonymous reviewers for their constructive feedback.
Appendix A Neuron-wise Analysis
We present additional examples illustrating how the infant model perform classification using visual concept neurons. Furthermore, we provide complete results of Age of Acquisition (AoA) to quantify the cognitive level of visual concepts for both in-vocabulary and out-of-vocabulary cases.
A.1 Neuron-based Classification Examples
A.2 Age-of-Acquisition Ratings
Age-of-Acquisition (AoA) ratings, defined by Kuperman et al. [28], estimate the age at which a person learns a word. These ratings were obtained via crowdsourcing using 30,121 English content words, organized into frequency-matched lists based on the SUBTLEX-US corpus [4]. Each list included calibrator and control words for validation. It strongly correlated with prior norms ($r=0.93$ with Cortese and Khanna [9], $r=0.86$ with the Bristol norms [38]), confirming their reliability for studying vocabulary development.
A.2.1 Detailed AoA Results
Here we present visual concepts that from Konkle object dataset[25] class, by applying neuron labeling, we found many visual concept neurons with corresponding class inside vision encoder’s hidden representation. For founded classes, we investigate their AoA values to prove the alignment between computational model and infant cognition.
Vocab (Col 1) AoA (Col 1) Vocab (Col 2) AoA (Col 2)
bike 2.9 abagel 4.79
stamp 2.94 umbrell 4.79
microwave 3.23 desk 5.00
pen 3.33 hat 5.11
knife 3.37 cookie 5.50
broom 3.43 stool 5.56
scissors 4.05 necklace 5.61
button 4.15 sofa 5.63
hairbrush 4.15 fan 5.68
pizza 4.26 chair 6.00
kayak 4.42 ball 6.21
bucket 4.5 sandwich 6.33
clock 4.5 pants 7.67
apple 4.67 socks (sock) 8.80
tricycle 4.7 bowl 8.90
camera 4.78
Vocab (Col 1) AoA (Col 1) Vocab (Col 2) AoA (Col 2)
sippycup (cup) 3.57 collar 6.56
toyrabbit (rabbit) 3.94 yarn 6.61
toyhorse (horse) 4.15 necktie 6.63
dresser 4.28 hanger 6.78
roadsign (sign) 4.32 binoculars 6.79
rug 4.61 telescope 6.95
doorknob 4.70 seashell 7.06
mask 4.80 golfball (golf) 7.16
dollhouse 4.86 dumbbell 7.56
muffins (muffin) 5.11 bathsuit (bathrobe) 7.90
tent 5.16 bowtie 7.94
hammer 5.42 rosary 8.21
frisbee 5.50 calculator 8.22
cushion 5.53 suitcase 8.22
watergun (gun) 5.58 trunk 8.30
ceilingfan (fan) 5.63 chessboard 8.37
helmet 5.71 compass 8.44
stapler 5.83 cupsaucer (saucer) 8.44
axe 6.11 lantern 8.55
speakers (speaker) 6.11 licenseplate (license) 8.70
lawnmower 6.11 pokercard (poker) 9.10
domino 6.17 keyboard 9.32
recordplayer 6.37 ringbinder (binder) 10.42
pitcher 6.42 powerstrip 12.01
grill 6.53
A.3 Further Clarification on $n$-Way Classification Results
In Figure 8, the infant model CVCL ( ) using “ours” has limited class coverage. As shown in the Venn diagram (Figure 5) and the coverage results (Figure 6), the “in-vocab” class in this setting is restricted. Consequently, the results include at most 31 classes. Therefore, in Figure 8, the rightmost point for “CVCL-ResNeXt50 (Ours)” corresponds to $n=31$ instead of $n=32$.
Appendix B Centered Kernel Alignment (CKA)
@@
For two sets of activations, $\mathbf{X}\in\mathbb{R}^{n\times p}$ and $\mathbf{Y}\in\mathbb{R}^{n\times q}$, from corresponding layers of two models, where $n$ is the number of examples and $p$ and $q$ are the feature dimensions (i.e., the number of neurons in each layer), the linear CKA is defined as:
@@
Appendix C Call for More Openly Available Infant Dataset
The success of infant computational models [43, 34, 33] demonstrates the research potential of infant datasets like SAYCam [39]. While current access platform, e.g., Databrary [^1] https://databrary.org/ (requiring institutional agreements) prioritize participant privacy, we identify an opportunity to expand access to inspire more research innovations.
We call for more openly available infant datasets, similar to Ego4D [19] and Epic Kitchen [10], while ensuring robust privacy safeguards (e.g., by blurring faces, removing other privacy-sensitive information, and muting any personally identifiable audio). Since modifying existing datasets for open access may be constrained by prior agreements, we encourage the development of new infant datasets with greater openness and sufficient privacy protection measures.